OpenStudy (anonymous):

What is the reciprocal of -2/3 -3 -3/2 2/3 3/2

OpenStudy (anonymous):

B

OpenStudy (anonymous):

medal plzz

OpenStudy (anonymous):

i dont think that even the right answer so...

OpenStudy (babynini):

the reciprocal means putting the bottom number on top, and the top number on bottom! so yes B is correct :)

OpenStudy (babynini):

\[\frac{ -2 }{ 3 }\] becomes \[\frac{ 3 }{ -2 }\]

OpenStudy (babynini):

Just flip the numbers.

OpenStudy (anonymous):

thxs can u help with another question or more lol please

OpenStudy (anonymous):

wait @Babynini their is no chocie of 3/-2

OpenStudy (babynini):

oops sorry haha I wrote it wrong. but -3/2 is the same as writing 3/-2

OpenStudy (anonymous):

2/3

iYuko (iyuko):

-2

OpenStudy (babynini):

ooh my gosh you guys. the answer is B!! :P

OpenStudy (babynini):

what are your other questions? :)

OpenStudy (anonymous):

ikr thxs @Babynini What is the solution to the equation if you solve for a? 2a – 3bc = −9

OpenStudy (er.mohd.amir):

B -3/2

OpenStudy (anonymous):

their are options here u go

OpenStudy (er.mohd.amir):

2a=3bc-9 a=1/2(3bc-9)

OpenStudy (anonymous):

What is the reciprocal of −1? (Points : 1) −1 0 1

iYuko (iyuko):

1

iYuko (iyuko):

-1*

OpenStudy (er.mohd.amir):

recipocal = -1

OpenStudy (anonymous):

@Er.Mohd.AMIR can u say it in like A B or C lol i did understand it like that lol

OpenStudy (anonymous):

@Er.Mohd.AMIR i didn't understand the second question the answer u gave me lol

OpenStudy (er.mohd.amir):

a=(1/2)(2bc-9) in terms of b and c if other eq give it.

OpenStudy (babynini):

the reciprocal of -1 is -1

OpenStudy (anonymous):

yea so the other one is what C lol

OpenStudy (anonymous):

im sorry i just cant get if u guys are saying C or what im really sorry

OpenStudy (babynini):

Sweetheart you didn't give us the options for the answers

OpenStudy (anonymous):

yea i did i swear i did lol go up a little if u can see all the comments on here and their not to far to :)

OpenStudy (babynini):

ooh I see! Sorry sorry, I thought you were asking for the reciprocal of -1 lool haha ok, let me try that again

OpenStudy (anonymous):

okay lol its alright

OpenStudy (babynini):

ok so you have 2a-3bc=-9 you want to start by getting a on one side by itself. what's a first step that could help you get there?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@Er.Mohd.AMIR u have any ideas of what this is ?

OpenStudy (babynini):

Isolate 2a

OpenStudy (anonymous):

kk

OpenStudy (er.mohd.amir):

2a=-9+3bc a=(1/2) (-9+3bc)

OpenStudy (babynini):

2a-3bc=-9 Add 3bc to both sides 21= -9+3bc

OpenStudy (babynini):

(that 1 is supposed to be an a)

OpenStudy (babynini):

then to get the a by itself divide both sides by 2

OpenStudy (babynini):

\[a=\frac{ -9 }{ 2 }+\frac{ 3bc }{ 2 }\]

OpenStudy (babynini):

and there's your answer!

OpenStudy (anonymous):

thxs lol help with more please u guys are the best

OpenStudy (babynini):

sure thing :)

OpenStudy (babynini):

put it on a new question thread though! I like to give people medalzzz xD

OpenStudy (anonymous):

thxs :*

OpenStudy (anonymous):

What is the slope of the line? thats the question

OpenStudy (er.mohd.amir):

please give equation

OpenStudy (anonymous):

here are the options −2 0

OpenStudy (babynini):

Ah! this one is fairly easy :) the slope of the line can be found just by looking at this. in equations were y=ax the slope is always a.